Probability
Showing 50 of 417 · 417 total
- 251.P(A B)=P(A B).Medium
- 252.P(A B)=16, P( A B)...Medium
- 253.Statement-1: P(E/A) P(A/E)P(E).Medium
- 254.X B(n,p), P(X=2)=P(X=3).Medium
- 255.A,B mutually exclusive, P(A)=3x+13, P(B)=1-x4.Medium
- 256.A, B, C independently hit a target with probabilities 34,12,58.Medium
- 257.Given two independent events, P(exactly one)=2649, P(none)=1549.Medium
- 258.Probability of a man hitting a target is 25.Easy
- 259.A multiple choice exam has 5 questions, each with 3 options (1 correct).Medium
- 260.An urn has 2 balls, each white or black.Medium
- 261.Six students including A and B stand in a row.Medium
- 262.A number n selected from 1,,1000.Medium
- 263.X,Y events with P(X Y)=P(X Y).Medium
- 264.Three numbers chosen without replacement from 1,,8.Medium
- 265.A,B,C pairwise independent, P(C)>0, P(A B C)=0.Medium
- 266.C D, P(D)0.Easy
- 267.5 independent Bernoulli trials with success probability p.Medium
- 268.Four numbers chosen at random (no replacement) from 1,,20.Medium
- 269.An urn has 9 balls: 3 red, 4 blue, 2 green.Medium
- 270.One ticket selected from 50 tickets numbered 00--49.Medium
- 271.In B(n,14), if P(at least one success)910, find n in terms of logs.Medium
- 272.A die is thrown; A: number >3; B: number <5.Easy
- 273.P(A)=14, P(A B)=12, P(B A)=23.Easy
- 274.Two aeroplanes bomb a target; P(I hits)=0.3, P(II hits)=0.2; II bombs only if I misses.Medium
- 275.A pair of dice thrown independently 3 times.Medium
- 276.A telephone enquiry system gets Poisson-distributed calls, mean 5 per 10 minutes.Medium
- 277.Three houses available; three persons each apply for one house without consulting others.Easy
- 278.P(A B)=16, P(A B)=14, P( A)=14.Medium
- 279.X has Poisson distribution with mean 2.Medium
- 280.Probability A speaks truth is 45, for B is 34.Medium
- 281.Mean and variance of a binomial distribution are 4 and 2.Medium
- 282.X has a given distribution.Medium
- 283.5 horses in a race.Easy
- 284.Events A,B,C mutually exclusive: P(A)=3x+13, P(B)=1-x4, P(C)=1-2x2.Medium
- 285.Mean and variance of a binomial X are 4 and 2.Medium
- 286.P(A B)=34, P(A B)=14, P( A)=23.Medium
- 287.A math problem given to A, B, C with probabilities of solving 12,13,14.Easy
- 288.A die is tossed 5 times, getting an odd number is a "success".Easy
- 289.A and B are events such that [P( A B ) = 3/4],[P( A B ) = 1/4,] [P( A ) = 2/3] then [P( A …Medium
- 290.A dice is tossed 5 times.Easy
- 291.Events A, B, C are mutually exclusive events such that [P( A ) = 3x + 1 3,] [P( B ) = 1 - …Medium
- 292.Five horses are in a race.Easy
- 293.The mean and variance of a random variable X having binomial distribution are 4 and 2 resp…Easy
- 294.The mean and the variance of a binomial distribution are 4 and 2 respectively.Easy
- 295.The probability that A speaks truth is [4 5,] while the probability for B is [3 4.] The pr…Easy
- 296.Three houses are available in a locality.Easy
- 297.A random variable X has Poisson distribution with mean 2.Easy
- 298.Let A and B two events such that [P( A B ) = 1 6,] [P( A B ) = 1 4] and [P( A ) = 1 4,] wh…Medium
- 299.At a telephone enquiry system the number of phone cells regarding relevant enquiry follow …Easy
- 300.Two aeroplanes [ I] and [ I] I[ bomb a target in succession.Medium